To create a breakfast beverage, pineapple juice in two concentrations, 20% and 70%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 20 gallons of 20% juice is used, how many gallons of the 70% juice must be used to obtain a 60% pineapple juice solution?
Answer by stanbon(75887)

To create a breakfast beverage, pineapple juice in two concentrations, 20% and 70%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 20 gallons of 20% juice is used, how many gallons of the 70% juice must be used to obtain a 60% pineapple juice solution?
--------------------------
Equation:
pine + pine = pine
0.20*20 + 0.70x = 0.60(20+x)
Multiply thru by 100 to get:
20*20 + 70x = 60*20 + 60x
10x = 40*20
x = 80 gallons (amount of 70% pineapple needed in the mixture)
=============================================================
Cheers,
Stan H.